# SMILES has to go : Representation of Molecules via Algebraic Data Types

### v1

@@ 3 Algebraic Data Types
For example, the Maybe type is an ADT that can encapsulate a computation that may or may not yield a value, signified by its constructors Nothing or Just a, respectively. This provides a type-safe way to handle operations that could fail, or denote an abscence of a value rather than using zero – a concept that in other languages might be handled by null references or exceptions, which may introduce errors.
In contrast to Haskell’s strong, static type system, languages like Python employ dynamic typing, which can lead to a different class of runtime errors. Large-scale machine learning applications often written in Python, such as those using the PyTorch framework, must contend with the possibility of type-related errors during execution. Whilst this is a well known advantage of using Haskell, it is not the focus of the paper.
The ADT in this article provides a minimal representation for cheminformaticians to use and extend. One desirable characteristic of this representation, is that it allows electrons to participate in more than one covalent or hydrogen bond, which is a consequence of using the representation set up by Dietz, which this article will refer to as the Dietz Representation [Dietz, 1995]. Ferrocene, an important organometallic that is difficult to faithfully represent in current formats now has a straightforward and readable representation. The Dietz format also allows for resonance structures (with multiple contributing graphs) to be exhibited [Dietz, 1995]. Indeed, implementing precisely this was proposed as an advancement of future molecular representations by [Krenn et al., 2022]. Furthermore, with the 3D coordinates of each atom, the ADT can be used to calculate the dihedral angles, bond angles and bond lengths, which are also unavailable in SMILES strings. The addition of these adds minimally to the space complexity, yet can be crucial for predicting the physical properties of drugs. Whilst not done here, the representation could be extended to implement single-atom tautomerism (which for example is more general than prototautemorism, representable with hydrogen in the inChi V1 framework) [Pospisil et al., 2003].
As a consequence of being an ADT that derives Show and Read instances, molecules can be leveraged as a transmission representation by directly serialising them to Haskell compatible syntax. These can act as storage files which can be read from, when a program requires said molecule.
@@ 3.1 Record Syntax
data Atom = AtomS { atomID :: Integer,
                    symbol :: AtomicSymbol }
          | AtomC { atomID :: Integer,
                    symbol :: AtomicSymbol,
                    coordinate :: Coordinate }
data AtomicSymbol = C | O | H
data Coordinate = Coordinate { x :: Double,
                               y :: Double,
                               z :: Double }
@@
Example 1: Updating the coordinate of an atom
setCoordinate :: Coordinate -> Atom -> Atom
setCoordinate newCoordinate atomWC@(AtomC _ _ _) = atomWC { coordinate = newCoordinate }
setCoordinate _ atom = atom
@@
Example 2: Retrieving the x-coordinate of an atom
getXCoordinate :: Atom -> Maybe Double
getXCoordinate (AtomC _ _ coordinate) = Just (x coordinate)
getXCoordinate _ = Nothing
@@ 3.2 The Molecule ADT
{verbnobox}
[] data Molecule = Molecule atoms :: [Atom], bonds :: M.Map (Integer, Integer) BondType deriving (Show, Read)
data Atom = Atom atomID :: Integer, atomicAttr :: ElementAttributes, coordinate :: Coordinate, shells :: Shells deriving (Show, Read)
data BondType = HydrogenBond | CovalentBond delocNum :: Integer, atomIDs :: Maybe [Integer] | IonicBond deriving (Eq, Read, Show)
data AtomicSymbol = O | H | N | C | B | Fe | F | Cl | S | Br | P | I deriving (Eq, Read, Show)
data ElementAttributes = ElementAttributes symbol :: AtomicSymbol, atomicNumber :: Integer, atomicWeight :: Double deriving (Eq, Read, Show)
data Coordinate = Coordinate x :: Double , y :: Double , z :: Double deriving (Show, Eq, Read)
Figure 5: The core data types for representing molecules in the proposed algebraic data type (ADT) system. The Molecule type consists of a list of Atoms and a map of BondTypes, which connect pairs of atoms identified by their unique Integer IDs. Each Atom contains an atomID, ElementAttributes (including the AtomicSymbol, atomicNumber, and atomicWeight), Coordinate (3D spatial position), and Shells (electronic configuration). The BondType is a sum type that can represent HydrogenBond, CovalentBond (with delocNum for the number of delocalized electrons and an optional list for which atoms the electrons are delocalised over), or IonicBond.
@@
The BondType data type is a sum type that captures the different types of chemical bonds that can exist between atoms. It includes HydrogenBond, CovalentBond, and IonicBond as possible variants. The CovalentBond constructor takes additional fields, delocNum and atomIDs, which represent the delocalization number and the optional ring structure associated with the covalent bond. This allows for the representation of delocalized electrons and cyclic structures within the molecule. This allows for the accurate modeling of the dynamic distribution of electrons across multiple bonds within a molecule, capturing the essence of the Dietz representation. By incorporating this information, the data type can effectively represent resonant structures and account for the delocalization of electrons, which is crucial for understanding the electronic properties and reactivity of molecules. The atomIDs field, of type Maybe [Integer], provides a means to represent the interconnected bonds within a cyclic structure. It is an optional field, indicated by the Maybe type constructor, allowing for the presence or absence of a list of atomIDs across which electrons are delocalised associated with the covalent bond. This list of integers enables for instance the representation of cyclic structures within the molecule. This design choice of using the delocNum and atomIDs fields in the CovalentBond constructor offers a flexible framework capable of accurately depicting complex molecules, such as organometallic compounds like ferrocene. It provides a foundation for computing fractional bond orders, which is essential for molecules with resonant bonding characteristics, such as ozone ($O_{3}$).
@@ 3.2.1 Example of a Molecule : Benzene
{verbnobox}
[] benzene :: Molecule benzene = Molecule atoms = [ Atom atomID = 1, atomicAttr = elementAttributes C, coordinate = Coordinate 0.0 1.3948 0.0, shells = carbon , Atom atomID = 2, atomicAttr = elementAttributes C, coordinate = Coordinate 1.2079 0.6974 0.0, shells = carbon , Atom atomID = 3, atomicAttr = elementAttributes C, coordinate = Coordinate 1.2079 (-0.6974) 0.0, shells = carbon , Atom atomID = 4, atomicAttr = elementAttributes C, coordinate = Coordinate 0.0 (-1.3948) 0.0, shells = carbon , Atom atomID = 5, atomicAttr = elementAttributes C, coordinate = Coordinate (-1.2079) (-0.6974) 0.0, shells = carbon , Atom atomID = 6, atomicAttr = elementAttributes C, coordinate = Coordinate (-1.2079) 0.6974 0.0, shells = carbon , Atom atomID = 7, atomicAttr = elementAttributes H, coordinate = Coordinate 0.0 2.4732 0.0, shells = hydrogen , Atom atomID = 8, atomicAttr = elementAttributes H, coordinate = Coordinate 2.1431 1.2366 0.0, shells = hydrogen , Atom atomID = 9, atomicAttr = elementAttributes H, coordinate = Coordinate 2.1431 (-1.2366) 0.0, shells = hydrogen , Atom atomID = 10, atomicAttr = elementAttributes H, coordinate = Coordinate 0.0 (-2.4732) 0.0, shells = hydrogen , Atom atomID = 11, atomicAttr = elementAttributes H, coordinate = Coordinate (-2.1431) (-1.2366) 0.0, shells = hydrogen , Atom atomID = 12, atomicAttr = elementAttributes H, coordinate = Coordinate (-2.1431) 1.2366 0.0, shells = hydrogen ]
, bonds = M.fromList \$ getSymmetricBonds [ ((1,2), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((1,6), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((1,7), CovalentBond delocNum = 2, atomIDs = Nothing) , ((2,3), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((2,8), CovalentBond delocNum = 2, atomIDs = Nothing) , ((3,4), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((3,9), CovalentBond delocNum = 2, atomIDs = Nothing) , ((4,5), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((4,10), CovalentBond delocNum = 2, atomIDs = Nothing) , ((5,6), CovalentBond delocNum = 6, atomIDs = Just [1,2,3,4,5,6]) , ((5,11), CovalentBond delocNum = 2, atomIDs = Nothing) , ((6,12), CovalentBond delocNum = 2, atomIDs = Nothing) ]
Figure 6: Representation of the benzene molecule using the proposed algebraic data type (ADT) system. The atoms field contains a list of 12 Atom values, representing the six carbon atoms and six hydrogen atoms. Each atom is specified with its unique atomID, atomicAttr (including the AtomicSymbol, atomicNumber, and atomicWeight), coordinate (3D spatial position), and shells (electronic configuration). The bonds field is represented using a Map from tuples of atom IDs to BondType values, capturing the delocalized nature of the electrons in the aromatic ring and the explicit representation of the ring structure.
@@
The bonds field of the benzene molecule is represented using a Map from tuples of atom IDs to BondType values. The bonds between the carbon atoms in the ring are represented as CovalentBond values, with a delocNum of 6, indicating that there are six delocalized electrons shared among the carbon atoms in the ring. This accurately captures the aromatic nature of benzene, where the electrons are evenly distributed across the ring. The atomIDs field of each covalent bond in the carbon ring is set to Just [1,2,3,4,5,6], explicitly specifying the cyclic structure formed by the carbon atoms.
The bonds between the carbon atoms and their respective hydrogen atoms are also represented as CovalentBond values, but with a delocNum of 2, indicating a typical single bond between carbon and hydrogen. The atomIDs field for these bonds is therefore set to Nothing, as they are not part of the aromatic ring.
@@ 3.3 The Orbital ADT
{verbnobox}
[] – Orbital types for each subshell data So = So deriving (Show, Eq, Read) data P = Px | Py | Pz deriving (Show, Eq, Read) data D = Dxy | Dyz | Dxz | Dx2y2 | Dz2 deriving (Show, Eq, Read) data F = Fxxx | Fxxy | Fxxz | Fxyy | Fxyz | Fxzz | Fzzz deriving (Show, Eq, Read)
– The Orbital has a type and a number of electrons. data Orbital subshellType = Orbital orbitalType :: subshellType , electronCount :: Int , orientation :: Maybe Coordinate deriving (Show, Eq, Read)
– Each SubShell consists of a list of Orbitals. – We ensure that all orbitals within a subshell are of the same type – by using the subshellType parameter. newtype SubShell subshellType = SubShell orbitals :: [Orbital subshellType] deriving (Show, Eq, Read)
– Each Shell has a principal quantum number and consists of a list of SubShells. data Shell = Shell principalQuantumNumber :: Int , sSubShell :: Maybe (SubShell So) , pSubShell :: Maybe (SubShell P) , dSubShell :: Maybe (SubShell D) , fSubShell :: Maybe (SubShell F) deriving (Show, Eq, Read)
– An Atom consists of a list of Shells. type Shells = [Shell]
Figure 7: Haskell code defining the algebraic data types (ADTs) for representing the electronic configuration of atoms. The So, P, D, and F data types represent the different subshell types, while the Orbital data type captures the properties of an individual orbital, including its subshell type, electron count, and spatial orientation. The SubShell newtype encapsulates a list of orbitals of the same subshell type, and the Shell data type represents a shell with its principal quantum number and the corresponding subshells.
@@
The use of a list to represent the shells implies that there could be a countably infinite number of shells, which is consistent with the theoretical framework of quantum mechanics. According to the theory, the number of shells is unlimited, and each shell can accommodate a specific number of electrons related to the shell’s principal quantum number.
@@ 3.3.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
{verbnobox}
[] hydrogen :: Shells hydrogen = [ Shell principalQuantumNumber = 1 , sSubShell = Just (SubShell [Orbital So 1 Nothing]) – 1s^1 , pSubShell = Nothing , dSubShell = Nothing , fSubShell = Nothing ]
carbon :: Shells carbon = [ Shell principalQuantumNumber = 1 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 1s^2 , pSubShell = Nothing , dSubShell = Nothing , fSubShell = Nothing , Shell principalQuantumNumber = 2 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 2s^2 , pSubShell = Just (SubShell [Orbital Px 1 (Just (Coordinate 1 0 0)), Orbital Py 1 (Just (Coordinate 0 1 0))]) – 2p^2 , dSubShell = Nothing , fSubShell = Nothing ]
oxygen :: Shells oxygen = [ Shell principalQuantumNumber = 1 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 1s^2 , pSubShell = Nothing , dSubShell = Nothing , fSubShell = Nothing , Shell principalQuantumNumber = 2 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) – 2s^2 , pSubShell = Just (SubShell [Orbital Px 2 (Just (Coordinate 1 0 0)), Orbital Py 1 (Just (Coordinate 0 1 0)), Orbital Pz 1 (Just (Coordinate 0 0 1))]) – 2p^4 , dSubShell = Nothing , fSubShell = Nothing ]
Figure 8: Examples of electronic configurations for hydrogen, carbon, and oxygen atoms using the algebraic data type representation in Haskell.

### v2

@@ 5.1 Haskell
The Maybe and Nothing types provide a type-safe way to handle operations that may fail, distinguishing between empty and zero values, a concept that in other languages might be handled by null references or exceptions, which often introduce program crashing errors. This distinction is later used in the orbital data type, to denote the absence of a SubShell.
Additionally, Haskell’s type-safe type system offers compile-time safety. This can completely eliminate a class of runtime errors commonly encountered in dynamically-typed languages like Python. Machine learning applications written with Python, or the PyTorch framework, must contend with the possibility of type-related errors during execution, even with type-annotation.
It should be noted that concepts such as Algebraic Data Types, or Functional Purity, are not unique to Haskell, therefore Haskell is not strictly necessary for the implementation.
Record syntax is now briefly discussed, before introducing the ADT itself.
Furthermore, we have dihedral and bond angles already. We have the meta characters as operations in haskell such as folds for a for loop and filter or takeWhile for a while loop.
@@ 5.2 Record Syntax
1data Atom = AtomS { atomID :: Integer,
2 symbol :: AtomicSymbol }
3 | AtomC { atomID :: Integer,
4 symbol :: AtomicSymbol,
5 coordinate :: Coordinate }
6
7data AtomicSymbol = C | O | H
8
9data Coordinate = Coordinate { x :: Double,
10 y :: Double,
11 z :: Double }
Listing 2: Updating the coordinate of an atom
@@
Example 1: Updating the x-coordinate of an atom
1updateXCoordinate :: Double -> Atom -> Atom
2updateXCoordinate newX (AtomC atomID symbol coordinate) =
3 AtomC atomID symbol coordinate { x = newX }
4updateXCoordinate _ atom = atom
Listing 3: Modifying only the x-coordinate of an atom
@@ 3 AtomC atomID symbol coordinate { x = newX }
Example 2: Extracting the x-coordinate of an atom
1getXCoordinate :: Atom -> Maybe Double
2getXCoordinate (AtomC _ _ coordinate) = Just (x coordinate)
3getXCoordinate _ = Nothing
Listing 4: Retrieving the x-coordinate of an atom
@@ 5.3 The Molecule ADT
1
2data Molecule = Molecule {
3 atoms :: [Atom],
4 bonds :: M.Map (Integer, Integer) BondType
5 } deriving (Show, Read)
6
7data Atom = Atom {
8 atomID :: Integer,
9 atomicAttr :: ElementAttributes,
10 coordinate :: Coordinate,
11 shells :: Shells
12 } deriving (Show, Read)
13
14data BondType = Bond {delocNum :: Integer, atomIDs :: Maybe [Integer]}
15 deriving (Eq, Read, Show)
16
17data AtomicSymbol = O | H | N | C | B | Fe | F | Cl | S | Br | P | I
18 deriving (Eq, Read, Show)
19
20data ElementAttributes = ElementAttributes
21 { symbol :: AtomicSymbol,
22 atomicNumber :: Integer,
23 atomicWeight :: Double
24 } deriving (Eq, Read, Show)
25
26type EquilibriumBondLength = Angstrom
27newtype Angstrom = Angstrom Double deriving (Read, Show, Eq)
Listing 5: The core data types for representing molecules in the proposed algebraic data type (ADT) system. The Molecule type consists of a list of Atoms and a map of BondTypes, which connect pairs of atoms identified by their unique Integer IDs. Each Atom contains an atomID, ElementAttributes (including the AtomicSymbol, atomicNumber, and atomicWeight), Coordinate (3D spatial position), and Shells (electronic configuration). The BondType is a sum type that can represent Bonds (with delocNum for the number of delocalized electrons and an optional list for which atoms the electrons are delocalised over).
@@
The bonds field of the Molecule data type is represented using a Map, with the keys being tuples of Integer values representing the atom IDs of the bonded atoms, and the values being of type BondType. This representation acts as an adjacency matrix and allows for the efficient lookup and manipulation of bonds within the molecule.
The BondType data type includes delocNum and atomIDs, which represent the delocalization number and the optional ring structure associated with the covalent bond. This allows for the representation of delocalized electrons and cyclic structures within the molecule. This allows for the accurate modeling of the dynamic distribution of electrons across multiple bonds within a molecule, capturing the essence of the Dietz representation.
By incorporating this information, the data type can effectively represent resonant structures and account for the delocalization of electrons, which is crucial for understanding the electronic properties and reactivity of molecules. The atomIDs field, of type Maybe [Integer], provides a means to represent the interconnected bonds within a cyclic structure. It is an optional field, indicated by the Maybe type constructor, allowing for the presence or absence of a list of atomIDs across which electrons are delocalised associated with the covalent bond. This list of integers enables for instance the representation of cyclic structures within the molecule. This design choice of using the delocNum and atomIDs fields in the Bond constructor offers a flexible framework capable of accurately depicting complex molecules, such as organometallic compounds like ferrocene. It provides a foundation for computing fractional bond orders, which is essential for molecules with resonant bonding characteristics, such as ozone ($O_{3}$).
ADTs are composed of simple, atomic data types combined using constructs such as products (tuples or records) and sums (variants or unions).
@@ 5.3.1 Example of a Molecule : Benzene
1
2benzene :: Molecule
3benzene = Molecule
4 { atoms = [ Atom i (elementAttributes e) (Coordinate x y z)
5 | (i, e, x, y) <- atomsData ],
6 bonds = M.fromList \$ getSymmetricBonds bondsData
7 }
8
9atomsData :: [(Int, Element, Double, Double, Double)]
10atomsData =
11 [ (1, C, 0.0, 1.3948, 0.0), (2, C, 1.2079, 0.6974, 0.0),
12 (3, C, 1.2079, -0.6974, 0.0), (4, C, 0.0, -1.3948, 0.0),
13 (5, C, -1.2079, -0.6974, 0.0), (6, C, -1.2079, 0.6974, 0.0),
14 (7, H, 0.0, 2.4732, 0.0), (8, H, 2.1431, 1.2366, 0.0),
15 (9, H, 2.1431, -1.2366, 0.0), (10, H, 0.0, -2.4732, 0.0),
16 (11, H, -2.1431, -1.2366, 0.0), (12, H, -2.1431, 1.2366, 0.0)
17 ]
18
19bondsData :: [((Int, Int), Bond)]
20bondsData =
21 -- 6 single CC bonds between neighboring carbon atoms
22 [ ((1,2), single), ((1,6), single), ((2,3), single),
23 ((3,4), single), ((4,5), single), ((5,6), single),
24
25 -- 2c2e bonds between carbon and hydrogen atoms
26 ((1,7), single), ((2,8), single), ((3,9), single),
27 ((4,10), single), ((5,11), single), ((6,12), single),
28
29 -- delocalised bonding in the benzene ring
30 ((1,2), aromatic), ((1,6), aromatic), ((2,3), aromatic),
31 ((3,4), aromatic), ((4,5), aromatic), ((5,6), aromatic) ]
32 where
33 single = Bond {delocNum = 2, atomIDs = Nothing}
34 aromatic = Bond {delocNum = 6, atomIDs = Just [1,2,3,4,5,6]}
Listing 6: A representation of the molecule benzene expressed in the Molecule data type, written as a .hs file, and using list comprehension to ensure readability, excluding the Shells currently.
@@
The bonds field of the benzene molecule is represented using a Map from tuples of atom IDs to BondType values. The bonds between the carbon atoms in the ring are represented as Bond values, with a delocNum of 6, indicating that there are six delocalized electrons shared among the carbon atoms in the ring. This accurately captures the aromatic nature of benzene, where the electrons are evenly distributed across the ring. The atomIDs field of each covalent bond in the carbon ring is set to Just [1,2,3,4,5,6], explicitly specifying the cyclic structure formed by the carbon atoms.
The bonds between the carbon atoms and their respective hydrogen atoms are also represented as Bond values, but with a delocNum of 2, indicating a typical single bond between carbon and hydrogen. The atomIDs field for these bonds is therefore set to Nothing, as they are not part of the aromatic ring.
@@ 5.4 The Orbital ADT
The use of a list to represent the shells implies that there could be an arbitrarily large finite number of shells (or infinite), which is consistent with the theoretical framework of quantum mechanics. According to the theory, the number of shells is unlimited, and each shell can accommodate a specific number of electrons related to the shell’s principal quantum number.
@@
1-- Orbital types for each subshell
2data So = So deriving (Show, Eq, Read)
3data P = Px | Py | Pz deriving (Show, Eq, Read)
4data D = Dxy | Dyz | Dxz | Dx2y2 | Dz2 deriving (Show, Eq, Read)
5data F = Fxxx | Fxxy | Fxxz | Fxyy | Fxyz | Fxzz | Fzzz deriving (Show, Eq, Read)
6
7-- The Orbital has a type and a number of electrons.
8data Orbital subshellType = Orbital {
9 orbitalType :: subshellType
10 , electronCount :: Int
11 , orientation :: Maybe Coordinate
12 } deriving (Show, Eq, Read)
13
14-- Each SubShell consists of a list of Orbitals.
15-- We ensure that all orbitals within a subshell are of the same type
16-- by using the subshellType parameter.
17newtype SubShell subshellType = SubShell {
18 orbitals :: [Orbital subshellType]
19 } deriving (Show, Eq, Read)
20
21-- Each Shell has a principal quantum number and consists of a list of SubShells.
22data Shell = Shell {
23 principalQuantumNumber :: Int
24 , sSubShell :: Maybe (SubShell So)
25 , pSubShell :: Maybe (SubShell P)
26 , dSubShell :: Maybe (SubShell D)
27 , fSubShell :: Maybe (SubShell F)
28 } deriving (Show, Eq, Read)
Listing 7: Haskell code defining the algebraic data types (ADTs) for representing the electronic configuration of atoms. The So, P, D, and F data types represent the different subshell types, while the Orbital data type captures the properties of an individual orbital, including its subshell type, electron count, and spatial orientation. The SubShell newtype encapsulates a list of orbitals of the same subshell type, and the Shell data type represents a shell with its principal quantum number and the corresponding subshells.
@@
. Furthermore, the code creates a data structure to model shells, subshells, and orbitals. The Shell data type represents a shell with a principal quantum number and optional subshells ($s$, $p$, $d$, $f$). Each subshell is represented as a list of orbitals. The Orbital data type represents an individual orbitals with its type, electron count, and orientation. The more general Shells type is defined as a list of Shell, representing the complete electronic structure of an atom.
@@ 5.4.1 Example of Orbitals : Hydrogen, Carbon & Oxygen
1hydrogen :: Shells
2hydrogen =
3 [ Shell
4 { principalQuantumNumber = 1
5 , sSubShell = Just (SubShell [Orbital So 1 Nothing]) -- 1s^1
6 , pSubShell = Nothing
7 , dSubShell = Nothing
8 , fSubShell = Nothing
9 }
10 ]
11
12carbon :: Shells
13carbon =
14 [ Shell
15 { principalQuantumNumber = 1
16 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 1s^2
17 , pSubShell = Nothing
18 , dSubShell = Nothing
19 , fSubShell = Nothing
20 }
21 , Shell
22 { principalQuantumNumber = 2
23 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 2s^2
24 , pSubShell = Just (SubShell [Orbital Px 1 (Just (Coordinate 1 0 0)),
25 Orbital Py 1 (Just (Coordinate 0 1 0))])
26 -- 2p^2
27 , dSubShell = Nothing
28 , fSubShell = Nothing
29 }
30 ]
31
32oxygen :: Shells
33oxygen =
34 [ Shell
35 { principalQuantumNumber = 1
36 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 1s^2
37 , pSubShell = Nothing
38 , dSubShell = Nothing
39 , fSubShell = Nothing
40 }
41 , Shell
42 { principalQuantumNumber = 2
43 , sSubShell = Just (SubShell [Orbital So 2 Nothing]) -- 2s^2
44 , pSubShell = Just (SubShell [Orbital Px 2 (Just (Coordinate 1 0 0)),
45 Orbital Py 1 (Just (Coordinate 0 1 0)),
46 Orbital Pz 1 (Just (Coordinate 0 0 1))])
47 -- 2p^4
48 , dSubShell = Nothing
49 , fSubShell = Nothing
50 }
51 ]
Listing 8: Examples of electronic configurations for hydrogen, carbon, and oxygen using the Orbital Algebraic Data Type representation in Haskell.